OpenStudy (kitkat16):

A pyramid has a square base with sides of length s. The height of the pyramid is equal to 1/2 of the length of a side on the base. Which formula represents the volume of the pyramid? V=1/2 s^3 V=1/6 s^3 V=1/3 s^3 V=3s^3 V=6s^3

OpenStudy (kitkat16):

this is confusing v=(1)(w)(h)/3 Would it be V=1/3 S^3???